BeautifulAngel:

Match the circle equations in general form with their corresponding equations in standard form. Tiles x2 + y2 − 4x + 12y − 20 = 0 (x − 6)2 + (y − 4)2 = 56 x2 + y2 + 6x − 8y − 10 = 0 (x − 2)2 + (y + 6)2 = 60 3x2 + 3y2 + 12x + 18y − 15 = 0 (x + 2)2 + (y + 3)2 = 18 5x2 + 5y2 − 10x + 20y − 30 = 0 (x + 1)2 + (y − 6)2 = 46 2x2 + 2y2 − 24x − 16y − 8 = 0 x2 + y2 + 2x − 12y − 9 = 0

BeautifulAngel:

@Hero can you help with this tutorial?

Hero:

Okay, so starting with the first one \(x^2 + y^2 − 4x + 12y − 20 = 0\) First begin by pair x terms together and y terms together like so: \(x^2 - 4x + y^2 + 12y - 20 = 0\)

Hero:

Next, add 20 to both sides: \(x^2 - 4x + y^2 + 12y = 20\)

Hero:

So now we have two binomial expressions on the left side \(x^2 - 4x\) and \(y^2 + 12y\) Notice that -4 and 12 play the role of the coefficient b in the expression \(ax^2 + bx + c\) so we can add \(\left(\dfrac{b}{2}\right)^2\) to both sides for each binomial expression like so:

Hero:

\(x^2 - 4x + (-2)^2 + y^2 + 12y + (6)^2 = 20 + (-2)^2 + (6)^2\)

Hero:

@BeautifulAngel are you following so far?

BeautifulAngel:

Yes, just writing all this down

Hero:

The previous step above is knowing as completing the square, so if you are not familiar with that, there are plenty videos on it on youtube. The previous step simplifies to \(x^2 - 4x + 4 + y^2 + 12x + 36 = 20 + 4 + 36\) or just \(x^2 - 4x + 4 + y^2 + 12x + 36 = 60\) You can write each trinomial square on the left side in the form \(\left(x + \dfrac{b}{2}\right)^2\) like so \((x - 2)^2 + (y + 6)^2 = 60\)

Hero:

Which should match with one of your listed pairings.

BeautifulAngel:

If I did solved (x−2)2+(y+6)2=60 would it give me the other pair as well?

Hero:

Yes, you can reverse the process to get the other pair.

Hero:

Or rather, I should say you can get the other pair by expanding \((x - 2)^2 + (y + 6)^2 = 60\)

BeautifulAngel:

Would it be a more difficult process?

Hero:

You should do whichever is most convenient for you. Either way is the same level of difficulty. I don't think one way is more convenient than the other. I prefer to build than deconstruct. Whichever method you choose is a matter of preference.

BeautifulAngel:

Ok. Thank you for helping @Hero :)

Hero:

You're welcome.